OpenStudy (anonymous):

I have to find the boiling point of 0.17 m NaCl solution using the mass percent and the van't Hoff factors given to me. What exactly do I do to answer this?

OpenStudy (anonymous):

My mass percent of NaCl: 1)0.0070 2)0.500 3)1.00 4)2.00

OpenStudy (anonymous):

van't Hoff factors: 1)1.94 2)1.87 3)1.84 4)1.83

OpenStudy (taramgrant0543664):

This is the normal formula used ΔT = i Kb m where ΔT is the change in temperature i is the van't Hoff factor Kb is the boiling point constant and m is the moles/kg of solute

OpenStudy (anonymous):

So, where does my mass percent play a role in this equation?

OpenStudy (ramirocruzo):

ΔTb = i Kbm Kb = 0.512 Since: NaCl → Na+(aq) + Cl-(aq): m(Na+) = 0.17m m(Cl-) = 0.17m m(NaCl(aq)) = 0.17 + 0.17 = 0.34m ΔTb = i 0.34 x 0.512 = i 0.17408 C For the cases you have given, Mass % = Mass of Solute/Mass of Solvent Mass of solvent= Mass % x Mass of Solution Assume, 1 ltr H2O i.e. 1000gm, So, Mass of NaCl =1000 x Mass % Then, Molality(m) = moles solute particles per kg of solution = 1000 x Mass % / Gram Molecular Mass of Nacl =1000 x Mass %/58 Put values & have answers............ :D